Part of $15,000 is invested at 8% annual interest, and the rest is invested at 6%. After one year, the accounts paid $1,000 in interest. How much was invested at the lower rate?
Answer by mananth(16949)

Part I 8.00% per annum
Part II 6.00% per annum

x+y=15000------------------------1
8.00%x+ 6.00%y=$1,000.00
Multiply by 100
8x+6y=$100,000.00--------2

Multiply (1) by -8
we get

-8x-8y=-120000.00

Add this to (2)

-2y=-$20,000.00

divide by -2

y=$10,000.00 investment at 6.00%
Balance $5,000.00 investment at 8.00%